# High-dimensional Gaussian and bootstrap approximations for robust means

### v3

@@ Appendix B Preparatory lemmas
We proceed by bounding the second summand on the right-hand side of the previous display (the argument for the first summand is analogous and hence skipped). To this end, consider first the case of
$32b_{2}^{2}\mathinner{\Bigl(\frac{3c-1}{c-1}\Bigr)}\mathinner{\Bigl(\frac{c+1}{c}\Bigr)}\varepsilon_{n}^{1-\frac{2}{m}}\leq\frac{b_{1}^{2}}{2},$ (C.3)
which by Lemma B.1 implies that $\min_{j=1,\ldots,d}\overline{\Sigma}_{\varepsilon_{n},j,j}\geq b_{1}^{2}/2$. Next, by the definition of $\overline{Y}_{n}$
@@
whence, noting also that $I_{n,1}$ and $\overline{I}_{n,3}$ are non-random, Lemma B.2 applied with $\mathfrak{b}_{1}^{2}=b_{1}^{2}/2$ implies the existence of a constant $C_{1}$ depending only on $b_{1},b_{2},c$ and $m$ such that
$\sup_{t\in\overline{\mathbb{R}}^{d}}\mathinner{\!\biggl\lvert\mathbb{P}\mathinner{\bigl(\overline{Y}_{n}\leq t\bigr)}-\mathbb{P}\mathinner{\biggl(\overline{Z}\leq t-\sqrt{n}\mathinner{(I_{n,1}+\overline{I}_{n,3})}\biggr)}\biggr\rvert}\leq C_{1}\mathinner{\biggl(\frac{\log^{5-\frac{2}{m}}(dn)}{n^{1-\frac{2}{m}}}\biggr)}^{\frac{1}{4}},$
@@
$\geq-C_{2}\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{1}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{m}}\biggr)}\sqrt{n\log(d)},$
@@
the final inequality following from Lemma B.3, (C.1), $\min_{j=1,\ldots,d}\overline{\Sigma}_{\varepsilon_{n},j,j}\geq b_{1}^{2}/2$, and $C_{2}$ being a constant depending on $b_{1}$, $b_{2}$$c$ and $m$ only. Thus, combining the previous two displays, there exists a constant $C_{3}$ depending on $b_{1},b_{2},c$ and $m$ only such that when (C.3) is satisfied one has
@@
$\sup_{t\in\overline{\mathbb{R}}^{d}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(\overline{Y}_{n}\leq t\bigr)}-\mathbb{P}\mathinner{\bigl(\overline{Z}\leq t\bigr)}\Bigr\rvert}\leq C_{3}\mathinner{\Biggl(\mathinner{\biggl[\frac{\log^{5-\frac{2}{m}}(dn)}{n^{1-\frac{2}{m}}}\biggr]}^{\frac{1}{4}}+\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{1}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{m}}\biggr]}\sqrt{n\log(d)}\Biggr)}.$
@@
Finally, by the Gaussian-to-Gaussian comparison inequality as stated in Proposition 2.1 of Chernozhuokov et al. (2022) and Lemma B.1, there exists a constant $C_{4}$, depending on $b_{1}$$b_{2}$ and $c$ only ($C_{4}$ changes value in the second inequality below), such that
@@
$\sup_{t\in\overline{\mathbb{R}}^{d}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(\overline{Z}\leq t\bigr)}-\mathbb{P}\mathinner{\bigl(Z\leq t\bigr)}\Bigr\rvert}$ $\leq C_{4}\mathinner{\Bigl(\log^{2}(d)\varepsilon_{n}^{1-\frac{2}{m}}\Bigr)}^{1/2}$
$\leq C_{4}\mathinner{\Biggl(\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{2}{m}}\biggr]}\Biggr)}^{1/2}.$
@@
The previous two displays yield that there exists a constant $C$ depending only on $b_{1}$$b_{2}$$c$ and $m$ such that when (C.3) is satisfied it holds that
@@
$\sup_{t\in\overline{\mathbb{R}}^{d}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(\overline{Y}_{n}\leq t\bigr)}-\mathbb{P}\mathinner{\bigl(Z\leq t\bigr)}\Bigr\rvert}\leq C\mathinner{\Biggl(\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{2}{m}}\biggr]}\Biggr)}^{1/2}$
@@
$+\ C\mathinner{\Biggl(\mathinner{\biggl[\frac{\log^{5-\frac{2}{m}}(dn)}{n^{1-\frac{2}{m}}}\biggr]}^{\frac{1}{4}}+\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{1}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{m}}\biggr]}\sqrt{n\log(d)}\Biggr)}.$
@@
This implies the claimed bound on $\rho_{n,W}$ in (6) by using that $4/n$ in (C) is dominated by, e.g., $\mathinner{\bigl[\frac{\log(dn)}{n}\bigr]}^{1-\frac{1}{m}}\sqrt{n\log(d)}$ (if necessary adjust $C$).
If, on the other hand, (C.3) is not satisfied then
@@
$32b_{2}^{2}\mathinner{\Bigl(\frac{3c-1}{c-1}\Bigr)}\mathinner{\Bigl(\frac{c+1}{c}\Bigr)}\varepsilon_{n}^{1-\frac{2}{m}}>\frac{b_{1}^{2}}{2}\quad\iff\quad\varepsilon_{n}>\mathinner{\biggl[\frac{b_{1}^{2}}{64b_{2}^{2}\mathinner{\bigl(\frac{3c-1}{c-1}\bigr)}\mathinner{\bigl(\frac{c+1}{c}\bigr)}}\biggr]}^{\frac{m}{m-2}}=\mathrel{\mathop{\ordinarycolon}}K(b_{1},b_{2},c,m),$
where $K=K(b_{1},b_{2},c,m)$ does not depend on $n$. Thus, by the definition of $\varepsilon_{n}$ (cf. also Footnote 4)
$\overline{\eta}_{n}>\frac{K}{2\lambda_{1,c}}\qquad\text{or}\qquad\frac{\log(dn)}{n}>\frac{K}{2f(c)},$
@@
where $f(c)=\frac{c}{3[1-\sqrt{2(c^{2}-1)}]}\vee c\mathinner{\Bigl(\sqrt{2\frac{c+1}{c-1}}+\frac{1}{3}\Bigr)}$. In either case, since $\rho_{n,W}\leq 1$ the bound in (6) remains valid by adjusting the constant $C$, if necessary.
@@ Appendix D Proofs for Section 3
In order to prove Theorem 3.1, we introduce an intermediate estimator $\hat{\Sigma}_{n}$ of $\Sigma$ that may be of independent interest. Its properties are established by an application of the union bound (over the entries of the covariance matrix) to the one-dimensional winsorized mean estimator in Theorem 5.1 of Kock and Preinerstorfer (2025). Subsequently, the properties of $\tilde{\Sigma}_{n}$ are established by showing it is sufficiently close to $\hat{\Sigma}_{n}$.
To define $\hat{\Sigma}_{n}$, recall that $Y_{i}=\frac{1}{\sqrt{2}}\mathinner{(X_{2i}-X_{2i-1})}$ and $\tilde{Y}_{i}=\frac{1}{\sqrt{2}}\mathinner{(\tilde{X}_{2i}-\tilde{X}_{2i-1})}$ for $i=1,\ldots,N=n/2$, cf. (8). Next, write $U_{i,j,k}=Y_{i,j}Y_{i,k}$ as well as $\tilde{U}_{i,j,k}=\tilde{Y}_{i,j}\tilde{Y}_{i,k}$ for $i=1,\ldots,n/2$ and $1\leq j,k\leq d$. For $\varepsilon_{n}^{\prime}$ as in (9), define $\hat{\alpha}_{j,k}=\tilde{U}^{*}_{\lceil\varepsilon_{n}^{\prime}N\rceil,j,k}$ as well as $\hat{\beta}_{j,k}=\tilde{U}^{*}_{\lceil(1-\varepsilon_{n}^{\prime})N\rceil,j,k}$. Finally, let $\hat{\Sigma}_{n}$ be the estimator with elements
$\hat{\Sigma}_{n,j,k}=\frac{1}{N}\sum_{i=1}^{N}\phi_{\hat{\alpha}_{j,k},\hat{\beta}_{j,k}}\mathinner{\bigl(\tilde{U}_{i,j,k}\bigr)},\qquad 1\leq j,k\leq d.$ (D.1)
Theorem D.1.
Fix $c\in(1,\infty)$, and let Assumption 2.1 be satisfied with $m>2$. If $\varepsilon_{n}^{\prime}\in(0,1/2)$, for $\varepsilon_{n}^{\prime}$ as in (9), then for a constant $C=C(b_{2},c,m)$
@@
$\mathbb{P}\mathinner{\Biggl(\max_{1\leq j,k\leq d}\mathinner{\!\bigl\lvert\hat{\Sigma}_{n,j,k}-\Sigma_{j,k}\bigr\rvert}>C\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{\frac{1}{2}-\frac{1}{m}}\biggr]}\Biggr)}\leq\frac{12}{n}.$ (D.2)
Remark D.1.
The estimator $\hat{\Sigma}_{n}$, although symmetric, need not be positive semi-definite (PSD). If $\hat{\Sigma}_{n}$ is not PSD (which can easily be checked), PSD can be enforced by projecting $\hat{\Sigma}_{n}$ onto the convex cone $\mathbf{S}_{+}^{d}$ of $d\times d$ symmetric PSD matrices in the maximum entry-wise distance. Thus, one replaces $\hat{\Sigma}_{n}$ by [^8] The convex set of minimizers is non-empty (and closed) by, e.g., Theorem 27.1 in Rockafellar (1997).
$\hat{\Sigma}_{n,\text{PSD}}\in\operatorname*{arg\,min}_{S\in\mathbf{S}_{+}^{d}}\max_{1\leq j,k\leq d}\mathinner{\!\bigl\lvert S_{j,k}-\hat{\Sigma}_{n,j,k}\bigr\rvert}$.
By the minimizing property of $\hat{\Sigma}_{n,\text{PSD}}$ and $\Sigma\in\mathbf{S}_{+}^{d}$, the triangle inequality implies that
$\max_{1\leq j,k\leq d}\mathinner{\!\bigl\lvert\hat{\Sigma}_{n,j,k,\text{PSD}}-\Sigma_{j,k}\bigr\rvert}$ $\leq\max_{1\leq j,k\leq d}\mathinner{\!\bigl\lvert\hat{\Sigma}_{n,j,k,\text{PSD}}-\hat{\Sigma}_{n,j,k}\bigr\rvert}+\max_{1\leq j,k\leq d}\mathinner{\!\bigl\lvert\hat{\Sigma}_{n,j,k}-\Sigma_{j,k}\bigr\rvert}$ (D.3)
$\leq 2\max_{1\leq j,k\leq d}\mathinner{\!\bigl\lvert\hat{\Sigma}_{n,j,k}-\Sigma_{j,k}\bigr\rvert}$,
so that $\hat{\Sigma}_{n,\text{PSD}}$ also satisfies (D.2) (with $C$ replaced by $2C$).
Proof of Theorem D.1.
We set up for an elementwise application of Theorem 5.1 in Kock and Preinerstorfer (2025). To this end, fix $1\leq j,k\leq d$. Note that $U_{i,j,k},\ i=1,\ldots,n/2$ is i.i.d. with $\mathbb{E}U_{1,j,k}=\Sigma_{j,k}$. Furthermore, at most $\overline{\eta}_{n}n=2\overline{\eta}_{n}N$ of the $\tilde{U}_{i,j,k}$ differ from $U_{i,j,k}$. Next, by the Cauchy-Schwarz inequality,
$\mathbb{E}\mathinner{\!\bigl\lvert U_{1,j,k}-\mathbb{E}U_{1,j,k}\bigr\rvert}^{m/2}$ $\leq 2^{m/2}\mathbb{E}\mathinner{\!\bigl\lvert U_{1,j,k}\bigr\rvert}^{m/2}=\mathbb{E}\mathinner{\!\bigl\lvert\mathinner{\bigl(X_{2,j}-X_{1,j}\bigr)}\mathinner{\bigl(X_{2,k}-X_{1,k}\bigr)}\bigr\rvert}^{m/2}$
$\leq\mathinner{\Bigl(\mathbb{E}\mathinner{\!\bigl\lvert X_{2,j}-X_{1,j}\bigr\rvert}^{m}\mathbb{E}\mathinner{\!\bigl\lvert X_{2,k}-X_{1,k}\bigr\rvert}^{m}\Bigr)}^{1/2}.$
Thus, since for all $j=1,\ldots,d$
$\mathbb{E}\mathinner{\!\bigl\lvert X_{2,j}-X_{1,j}\bigr\rvert}^{m}=\mathbb{E}\mathinner{\!\bigl\lvert(X_{2,j}-\mu_{j})-(X_{1,j}-\mu_{j})\bigr\rvert}^{m}\leq 2^{m}\sigma_{m}^{m},$
the two previous displays imply that
$\mathinner{\Bigl(\mathbb{E}\mathinner{\!\bigl\lvert U_{1,j,k}-\mathbb{E}U_{1,j,k}\bigr\rvert}^{m/2}\Bigr)}^{2/m}\leq 4\sigma_{m}^{2}.$ (D.4)
@@
$\mathbb{P}\mathinner{\Biggl(\mathinner{\!\bigl\lvert\hat{\Sigma}_{n,j,k}-\Sigma_{j,k}\bigr\rvert}>C\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{\frac{1}{2}-\frac{1}{m}}\biggr]}\Biggr)}\leq\frac{12}{d^{2}n},$
where $C$ depends on $b_{2},c$, and $m$ only and we have inserted $N=n/2$. (D.2) now follows by the union bound. ∎
Remark D.2.
If $m\geq 4$, one can use $\varepsilon_{n}^{\prime}$ as in (12) and instead appeal to Theorem 3.1 in Kock and Preinerstorfer (2025) [this value of $\varepsilon_{n}^{\prime}$ being implied by the same choices as right after (D.4)] to show that the resulting estimator $\check{\Sigma}_{n}$ satisfies
@@
$\mathbb{P}\mathinner{\Biggl(\max_{1\leq j,k\leq d}\mathinner{\!\bigl\lvert\check{\Sigma}_{n,j,k}-\Sigma_{j,k}\bigr\rvert}>C\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{\frac{1}{2}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{1-\frac{2}{m}}\biggr]}\Biggr)}\leq\frac{12}{n},$
and thus has a better dependence on $n$ and $d$ than $\hat{\Sigma}_{n}$.
Theorem D.2.
Fix $c\in(1,\sqrt{1.5})$, let Assumption 2.1 be satisfied with $m>2$. Let $n>12$. If $\varepsilon_{n},\varepsilon_{n}^{\prime}\in(0,1/2)$, with $\varepsilon_{n}$ as in (5) and $\varepsilon_{n}^{\prime}$ as in (9), then for $\hat{Z}\sim\mathsf{N}_{d}(0,\hat{\Sigma}_{n,\mathrm{PSD}})$ conditionally on $\hat{\Sigma}_{n,\mathrm{PSD}}$ as in Remark D.1, cf. also (D.1), positive semi-definite, it holds with probability at least $1-\frac{12}{n}$ that
$\hat{\rho}_{n,W}\mathrel{\mathop{\ordinarycolon}}$ $=\sup_{H\in\mathcal{H}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(S_{n,W}\in H\bigr)}-\mathbb{P}\mathinner{\bigl(\hat{Z}\in H\mid\tilde{X}_{1},\ldots,\tilde{X}_{n}\bigr)}\Bigr\rvert}$
@@
$\leq\mathfrak{A}_{n}+C\mathinner{\Biggl(\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl(\frac{\log(dn)}{n}\Bigr)}^{\frac{1}{2}-\frac{1}{m}}\biggr]}\Biggr)}^{1/2}$
where $\mathfrak{A}_{n}$ is the upper bound on $\rho_{n,W}$ in (6) of Theorem 2.1 and $C$ is a constant depending only on $b_{1},b_{2},c$ and $m$.
In particular, $\hat{\rho}_{n,W}\to 0$ in probability if $\sqrt{n\log(d)}\overline{\eta}_{n}^{1-\frac{1}{m}}\to 0$ and $\log(d)/n^{\frac{m-2}{5m-2}}\to 0$.
Proof of Theorem D.2.
By the triangle inequality $\hat{\rho}_{n,W}$ is bounded from above by the sum of
where we recall that $Z\sim\mathsf{N}_{d}(0,\Sigma)$, and

### v4

@@ Appendix B Preparatory lemmas
We proceed by bounding the second summand on the right-hand side of the previous display (the argument for the first summand is analogous and hence skipped). To this end, note that (C.1) and (C.2) imply that $\mathfrak{T}_{n}^{(2)}\leq b_{1}^{2}/2$ which then together with Lemma B.1 implies that $\min_{j=1,\ldots,d}\overline{\Sigma}_{\varepsilon_{n},j,j}\geq b_{1}^{2}/2$. Next, by the definition of $\overline{Y}_{n}$
@@
whence, noting also that $I_{n,1}$ and $\overline{I}_{n,3}$ are non-random, Lemma B.2 applied with $\mathfrak{b}_{1}^{2}=b_{1}^{2}/2$ implies the existence of a constant $C_{3}$ depending only on $b_{1},b_{2},\lambda_{1},\lambda_{2}$, and $m$ such that
$\sup_{t\in\overline{\mathbb{R}}^{d}}\mathinner{\!\biggl\lvert\mathbb{P}\mathinner{\bigl(\overline{Y}_{n}\leq t\bigr)}-\mathbb{P}\mathinner{\biggl(\overline{Z}\leq t-\sqrt{n}\mathinner{(I_{n,1}+\overline{I}_{n,3})}\biggr)}\biggr\rvert}\leq C_{3}\mathinner{\biggl(\frac{\log^{5-\frac{2}{m}}(dn)}{n^{1-\frac{2}{m}}}\biggr)}^{\frac{1}{4}},$
@@
$\geq-C_{4}\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{1}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{m}}\biggr)}\sqrt{n\log(d)},$
@@
the final inequality following from Lemma B.3, (C.3), $\min_{j=1,\ldots,d}\overline{\Sigma}_{\varepsilon_{n},j,j}\geq b_{1}^{2}/2$, and $C_{4}$ being a constant depending on $b_{1}$, $b_{2}$$\lambda_{1},\lambda_{2}$, and $m$ only. Thus, combining the previous two displays, there exists a constant $C_{5}$ depending on $b_{1},b_{2},\lambda_{1},\lambda_{2}$, and $m$ only such that when (C.2) is satisfied one has
@@
$\sup_{t\in\overline{\mathbb{R}}^{d}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(\overline{Y}_{n}\leq t\bigr)}-\mathbb{P}\mathinner{\bigl(\overline{Z}\leq t\bigr)}\Bigr\rvert}\leq C_{5}\mathinner{\Biggl(\mathinner{\biggl[\frac{\log^{5-\frac{2}{m}}(dn)}{n^{1-\frac{2}{m}}}\biggr]}^{\frac{1}{4}}+\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{1}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{m}}\biggr]}\sqrt{n\log(d)}\Biggr)}.$
@@
Finally, by the Gaussian-to-Gaussian comparison inequality as stated in Proposition 2.1 of Chernozhukov et al. (2022) and Lemma B.1, there exists a constant $C_{6}$, depending on $b_{1}$$b_{2}$ $\lambda_{1}$$\lambda_{2}$, and $m$ only ($C_{6}$ changes value in the second inequality below), such that
@@
$\sup_{t\in\overline{\mathbb{R}}^{d}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(\overline{Z}\leq t\bigr)}-\mathbb{P}\mathinner{\bigl(Z\leq t\bigr)}\Bigr\rvert}$ $\leq C_{6}\mathinner{\Bigl(\log^{2}(d)\varepsilon_{n}^{1-\frac{2}{m}}\Bigr)}^{1/2}$
$\leq C_{6}\mathinner{\Biggl(\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{2}{m}}\biggr]}\Biggr)}^{1/2}.$
@@
The previous two displays yield that there exists a constant $C_{7}$ depending only on $b_{1}$$b_{2}$$\lambda_{1},\lambda_{2}$, and $m$ such that when (C.2) is satisfied it holds that
@@
$\sup_{t\in\overline{\mathbb{R}}^{d}}\mathinner{\!\Bigl\lvert\mathbb{P}\mathinner{\bigl(\overline{Y}_{n}\leq t\bigr)}-\mathbb{P}\mathinner{\bigl(Z\leq t\bigr)}\Bigr\rvert}\leq C_{7}\mathinner{\Biggl(\log^{2}(d)\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{2}{m}}\biggr]}\Biggr)}^{1/2}$
@@
$+\ C_{7}\mathinner{\Biggl(\mathinner{\biggl[\frac{\log^{5-\frac{2}{m}}(dn)}{n^{1-\frac{2}{m}}}\biggr]}^{\frac{1}{4}}+\mathinner{\biggl[\overline{\eta}_{n}^{1-\frac{1}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{m}}\biggr]}\sqrt{n\log(d)}\Biggr)}.$
@@
This implies the claimed bound on $\rho_{n,W}$ in (5) by using that $4/n$ in (C) is dominated by, e.g., $\mathinner{\bigl[\frac{\log(dn)}{n}\bigr]}^{1-\frac{1}{m}}\sqrt{n\log(d)}$ (if necessary adjust $C_{7}$).
If, on the other hand, (C.2) is not satisfied, then
@@
$C_{1}\varepsilon_{n}^{1-\frac{2}{m}}\geq\frac{(b_{1}^{2}\wedge 1)}{2}\quad\iff\quad\varepsilon_{n}\geq\mathinner{\biggl[\frac{(b_{1}^{2}\wedge 1)}{2C_{1}}\biggr]}^{\frac{m}{m-2}}=\mathrel{\mathop{\ordinarycolon}}K(b_{1},b_{2},\lambda_{2},m),$
where $K=K(b_{1},b_{2},c,m)$ does not depend on $n$. Thus, by the definition of $\varepsilon_{n}$,
$\overline{\eta}_{n}\geq\frac{K}{2\lambda_{1}}\qquad\text{or}\qquad\frac{\log(dn)}{n}\geq\frac{K}{2\lambda_{2}},$
@@
In either case, since $\rho_{n,W}\leq 1$ the bound in (5) remains valid by adjusting the constant $C$ there, if necessary.
@@ Appendix D Proofs for Section 3
We start with the proof of Theorem 3.1. Before we give the proof, we first define $c_{1,n}^{\prime}$ and $c_{2,n}^{\prime}$ that show up in the proof: Recall the definition of $\varepsilon_{n}^{\prime}$ in (7). For $\lambda_{1}^{\prime}\in(1,\infty)$, define $A_{+}^{\prime}=1-\lambda_{1}^{\prime-1}\mathds{1}\mathinner{\{\overline{\eta}_{n}>0\}}$$A_{-}^{\prime}=1+\lambda_{1}^{\prime-1}\mathds{1}\mathinner{\{\overline{\eta}_{n}>0\}}$, and
$c_{1,n}^{\prime}\mathrel{\mathop{\ordinarycolon}}=-A_{+}^{\prime}W_{0}\mathinner{\bigl(-e^{-(\frac{\log(d^{2}n)}{\varepsilon_{n}^{\prime}n}+A_{+}^{\prime})/A_{+}^{\prime}}\bigr)}\in(0,A_{+}^{\prime})$ (D.1)
as well as
$c_{2,n}^{\prime}\mathrel{\mathop{\ordinarycolon}}=-A_{-}^{\prime}W_{-1}\mathinner{\bigl(-e^{-(\frac{\log(d^{2}n)}{\varepsilon_{n}^{\prime}n}+A_{-}^{\prime})/A_{-}^{\prime}}\bigr)}\in(A_{-}^{\prime},\infty);$ (D.2)
where we recall that $W_{0}$ and $W_{-1}$ denote the principal and lower branch, respectively, of Lambert’s $W$ function (cf., e.g., Corless et al. (1996)).
Proof of Theorem 3.1.
Fix $1\leq j,k\leq d$ and note that
$\tilde{\Sigma}_{n,j,k}$ $=\frac{1}{n}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\hat{a}_{j},\hat{b}_{j}}(\tilde{X}_{i,j})-\mu_{j}-(\tilde{\mu}_{n,j}-\mu_{j})\bigr]}\mathinner{\bigl[\phi_{\hat{a}_{k},\hat{b}_{k}}(\tilde{X}_{i,k})-\mu_{k}-(\tilde{\mu}_{n,k}-\mu_{k})\bigr]}$
$=\frac{1}{n}\sum_{i=1}^{n}\mathinner{\bigl[\phi_{\hat{a}_{j},\hat{b}_{j}}(\tilde{X}_{i,j})-\mu_{j}\bigr]}\mathinner{\bigl[\phi_{\hat{a}_{k},\hat{b}_{k}}(\tilde{X}_{i,k})-\mu_{k}\bigr]}-(\tilde{\mu}_{n,j}-\mu_{j})(\tilde{\mu}_{n,k}-\mu_{k})$
$=\frac{1}{n}\sum_{i=1}^{n}\phi_{\check{a}_{j},\check{b}_{j}}(\tilde{Y}_{i,j})\phi_{\check{a}_{k},\check{b}_{k}}(\tilde{Y}_{i,k})-(\tilde{\mu}_{n,j}-\mu_{j})(\tilde{\mu}_{n,k}-\mu_{k}),$
where $Y_{i,j}=X_{i,j}-\mu_{j}$$\tilde{Y}_{i,j}=\tilde{X}_{i,j}-\mu_{j}$$\check{a}_{j}=\tilde{Y}_{\lceil\varepsilon_{n}^{\prime}n\rceil,j}^{*}=\tilde{X}_{\lceil\varepsilon_{n}^{\prime}n\rceil,j}^{*}-\mu_{j}=\hat{a}_{j}-\mu_{j}$, and $\check{b}_{j}=\tilde{Y}_{\lceil(1-\varepsilon_{n}^{\prime})n\rceil,j}^{*}=\tilde{X}_{\lceil(1-\varepsilon_{n}^{\prime})n\rceil,j}^{*}-\mu_{j}=\hat{b}_{j}-\mu_{j}$.
By Theorem 2.1 in Kock and Preinerstorfer (2025) applied with $\delta=6/(d^{2}n)$ $\lambda_{i}=\lambda_{i}^{\prime}$ for $i=1,2$, there exists a constant $C_{1}$ depending only on $b_{2},\lambda_{1}^{\prime},\lambda_{2}^{\prime}$, and $m$, such that with probability at least $1-12/(d^{2}n)$
@@
$|\tilde{\mu}_{n,j}-\mu_{j}||\tilde{\mu}_{n,k}-\mu_{k}|$ $\leq C_{1}\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{1}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{m\wedge 2}}\biggr)}^{2}$
$\leq C_{1}\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{(m/2)\wedge 2}}\biggr)},$
the second estimate using that $\overline{\eta}_{n}$ and $\frac{\log(dn)}{n}$ are both strictly smaller than 1 by (9) (and potentially adjusting $C_{1}$).
@@
$\mathinner{\!\Bigl\lvert\frac{1}{n}\sum_{i=1}^{n}\phi_{\check{\alpha}_{j,k},\check{\beta}_{j,k}}(\tilde{U}_{i,j,k})-\Sigma_{j,k}\Bigr\rvert}\leq C_{2}\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{(m/2)\wedge 2}}\biggr)}.$
Thus, defining
$\mathfrak{S}_{n,j,k}=\mathrel{\mathop{\ordinarycolon}}\frac{1}{n}\sum_{i=1}^{n}\phi_{\check{a}_{j},\check{b}_{j}}(\tilde{Y}_{i,j})\phi_{\check{a}_{k},\check{b}_{k}}(\tilde{Y}_{i,k})-\frac{1}{n}\sum_{i=1}^{n}\phi_{\check{\alpha}_{j,k},\check{\beta}_{j,k}}(\tilde{U}_{i,j,k}),$
in order to prove the theorem, it suffices (by the union bound over $d^{2}$) to show that there exists a constant $C_{3}$ only on $b_{2},\lambda_{1}^{\prime},\lambda_{2}^{\prime}$, and $m$ such that with probability at least $1-6/(d^{2}n)$
@@
$\mathinner{\!\bigl\lvert\mathfrak{S}_{n,j,k}\bigr\rvert}\leq C_{3}\mathinner{\biggl(\overline{\eta}_{n}^{1-\frac{2}{m}}+\mathinner{\Bigl[\frac{\log(dn)}{n}\Bigr]}^{1-\frac{1}{(m/2)\wedge 2}}\biggr)}$ (D.3)
To establish (D.3), note that with